OpenStudy (anonymous):

Simplify the Equation Below, show steps and teach me how to do it. Thanks :)

OpenStudy (anonymous):

\[\sqrt{15}(3+\sqrt{5})\]

OpenStudy (anonymous):

Alright, you ready?

OpenStudy (anonymous):

yes mr. teacher

OpenStudy (anonymous):

Mentor/Tutor =0)

OpenStudy (anonymous):

lol ok mr. M

OpenStudy (anonymous):

Future professor.

OpenStudy (anonymous):

Alrighty. Do you know how to use the distributive property?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok well I cannot get it to copy and paste on here. Weird. Ok let's start by distributing using F.O.I.L. sqrt(15) X 3 = 3sqrt(15) + sqrt(15) X sqrt(5) = sqrt(75) So we have: 3sqrt(5) + sqrt(75) With me so far?

OpenStudy (jdoe0001):

\(\bf \sqrt{15}(3+\sqrt{5})\implies \sqrt{15}\cdot 3+\sqrt{15}\cdot \sqrt{5}\implies 3\sqrt{15}+\sqrt{15\cdot 5} \\ \quad \\ 3\sqrt{15}+\sqrt{75}\quad recall\to {\color{brown}{ 75\to 5\cdot 5\cdot 3\to 5^2\cdot 3}}\quad thus \\ \quad \\ 3\sqrt{15}+\sqrt{75}\implies 3\sqrt{15}+\sqrt{5^2\cdot 3}\implies 3\sqrt{15}+\large \sqrt[{\color{blue}{ 2}}]{5^{\color{blue}{ 2}}\cdot 3} \\ \quad \\ 3\sqrt{15}+5\sqrt{3}\)

OpenStudy (anonymous):

Im with you Professor Cosmic

OpenStudy (anonymous):

Actually refer to @jdoe0001's explanation as it is much more clear and precise!

OpenStudy (anonymous):

Let me know if you have any questions about how he came to 3sqrt(15) + 5sqrt(3)